# EnergyMoGen: Compositional Human Motion Generation with Energy-Based Diffusion Model in Latent Space

### v1

@@ 1 Introduction
In this work, we are interested in the motion latent diffusion model [52, 64], which is another important formulation of diffusion models operating in continuous latent space. However, such models are not feasible for composing motion through algorithms proposed in skeleton-based diffusion models. As latent diffusion models map motions of varying lengths into a fixed number of tokens, the following issues naturally arise: i) Using the fixed number of tokens to represent motions can not support per-frame composition; ii) Lack of explicit correspondence between the latent feature and skeleton.
@@
Taking these innovations together, EnergyMoGen becomes a general and flexible framework for compositional motion generation (our approach can easily generalize to skeleton-based diffusion models, and the results are provided in Section B of the appendix). Empirically, we show impressive results on three datasets across three tasks, i.e., HumanML3D [19] and KIT-ML [49] for text-to-motion generation task (§$\S$ 5.3), MTT [47] for compositional motion generation and multi-concept motion generation tasks (§$\S$ 5.4). We conduct comprehensive experiments for in-depth dissection (§$\S$ 5.5) to reveal the effectiveness of EnergyMoGen. Additionally, we introduce a CompML dataset (§$\S$ 5.1) comprising 5000 compositionally generated motion sequences with textual descriptions. We show that the model’s performance can be improved via training on composed motions (§$\S$ 5.5). We hope these experiments and conclusions will contribute to future developments.
@@ 2 Preliminaries: Energy-Based Models
In contrast to explicitly modeling the data distribution, Energy-based models (EBMs) [34, 74] implicitly define the underlying probability density of the distribution through an energy function. For a data sample ${\bm{X}}$, the probability density can be defined via the Boltzmann distribution:
@@
In addition, the modern Hopfield networks [50, 25] interpret the self-attention as EBMs. [50] introduces an energy function $E({\bm{X}},\xi)=-\alpha^{-1}\log\sum^{N}_{i=1}\exp(\alpha({\bm{X}}^{\top}\xi)_{i})+\frac{1}{2}\xi^{\top}\xi$ and a update rule $\xi_{new}=\xi-\gamma(\xi-{\bm{X}}\mathrm{softmax}(\alpha{\bm{X}}^{\top}\xi))$, where the data sample ${\bm{X}}$ consists of $N$ stored patterns (key) ${\bm{x}}_{i}\in\mathbb{R}^{d}$, $\xi\in\mathbb{R}^{d}$ denotes the state vector (query), and $\gamma$ is the update step size. By linearly projecting ${\bm{X}}$ and $\xi$ into Query (${\bm{Q}}$), Key(${\bm{K}}$), and Value(${\bm{V}}$), the attention mechanism can be understood as the energy update ${\bm{Q}}_{new}=\mathrm{softmax}(\alpha{\bm{Q}}{\bm{K}}){\bm{V}}$, when $\gamma=1$ and $\alpha=1/\sqrt{d}$. On this basis, Park et al. [43] extends the theory to cross-attention to handle several tasks, e.g., image editing and inpainting.
@@ 3.1 EnergyMoGen for Motion Generation
We build the architecture of EnergyMoGen upon motion latent diffusion model [64], which comprises two transformer-based modules: Motion Variational Autoencoder and Motion Latent Diffusion (in Figure 2 (a), (b)).
@@
Building upon the architecture in MLD [64], we propose to employ cross-attention to incorporate text embeddings and motion latent features. In light of recent endorsers on energy-based attention mechanism [50, 25], we interpret the cross-attention as an energy-based operation to facilitate multi-concept motion generation. Specifically, ${\bm{K}}$ and ${\bm{V}}$ in cross attention is computed as ${\bm{K}}={\bm{c}}W_{\bm{K}}$ and ${\bm{V}}={\bm{c}}W_{\bm{V}}$, where $W_{\bm{K}},W_{\bm{V}}\in\mathbb{R}^{d\times d}$. To alleviate the impact of misalignment in multi-concept text embeddings, we focus on refining ${\bm{c}}$ through adaptive gradient descent (AGD) based on MAP estimation with $p({\bm{K}}|{\bm{Q}})\propto p({\bm{Q}}|{\bm{K}})p({\bm{K}})$. We introduce two energy functions, $E({\bm{K}})$ and $E({\bm{Q}}|{\bm{K}})$, done in the prior work [43]. Then, the gradient of log distribution can be written as:
@@ 4 Related Works
Compositional Generation. Compositional generation has emerged as a powerful paradigm for creating content (e.g., images and motions) that combines multiple components. Our work is related to certain works in compositional image generation. Some approaches aim to develop cross-attention merging algorithms [12, 43], training strategies [35, 5, 26], and classifier guidance [54, 15], to facilitate this task. Another group of works proposes to use EBMs to compose pre-trained generative models [10, 7, 37, 38, 43, 8]. For example, Du et al. [8] proposed to combine independent trained energy models with several logical operators. Nie et al. [42] and Du et al. [11] introduced Ordinary Different Equation (ODE) and Markov Chain Monte Carlo (MCMC) to optimize the sampling process, respectively. In the field of motion generation, TEACH [2] and SINC [3] are pioneering works for spatial/temporal action composition based on VAE. Subsequent solutions [53, 30, 62] manipulate diffusion models to achieve more natural motion and controllability. However, they modify the reverse process in the skeleton space, which is infeasible for adapting to latent space.
@@ 5 Experiments
We evaluate our approach on three tasks following prior studies [19, 47]: text-to-motion generation (§$\S$ 5.3), compositional motion generation (§$\S$ 5.4), and multi-concept motion generation (§$\S$ 5.4). Furthermore, we offer an in-depth dissection in §$\S$ 5.5. Please kindly note that our approach can be easily generalized to skeleton-based methods. More information on datasets, evaluation metrics, training details, and the results of skeleton-based diffusion models, can be found in the appendix.
@@ 5.5 In-Depth Dissection of EnergyMoGen
Q1: Does the adaptive gradient descent alleviate the issue of text inconsistency in multi-concept motion generation? Yes. As shown in Table 3, our approach with adaptive updating of the textual embedding aligns better with textual descriptions. It improves the R-Precision by 1.3% and 0.9% for Top1 and Top3 accuracy, respectively. Meanwhile, we achieve better FID and comparable results on the TMR-Score and transition distance. The ablation study of $[\gamma_{attn},\gamma_{reg}]$ and qualitative comparisons are provided in Section E of the appendix.
Q2: What roles do latent-aware and semantic-aware EBMs play in motion composition? To answer this question, we conduct ablative experiments on MTT [47], results are shown in Table 3. One can figure out that the semantic-aware EBM significantly enhances the text-motion consistency, but it suffers from foot slides and motion jitter. Although the latent-aware method performs poorly on R-Precision and TMR Score, it clearly decreases the transition distance. We notice that Synergistic Energy Fusion (SEF) effectively alleviates these issues by leveraging the complementary strengths of the two spectrums. We present visual results and ablation studies on $\lambda_{s}$, $\lambda_{l}$, and $\lambda_{m}$ in Section C of the appendix.
@@ 6 Conclusion
  • Section F: More visual results of energy distributions.
  • Section G: More details on datasets and evaluation metrics.
@@ Appendix A Implementation Details
For the skeleton-based approach, we use an 8-layer transformer with a dimension of 512. We follow [69] to train the model using Adam optimizer with a batch size of 1024. We train 8000 epochs in total and employ the CosineAnnealing learning policy with the learning rate from 0.0002 to 0.00002.
@@ B.1 Text-to-Motion Generation
We conduct experiments on HumanML3D [19] to evaluate the performance of text-to-motion generation. We use evaluation models from Guo et al. [19] and use the same metrics. The training details of skeleton-based EnergyMoGen are provided in Section A. Experimental results are shown in Table 4. Our approach outperforms current state-of-the-art skeleton-based methods, i.e., ReMoDiffuse [70] and FineMoGen [71] on R-Precision, Diversity, and MM-Dist, while achieves comparable results on FID and MModality.
@@ B.3 Multi-Concept Motion Generation
Table 6: Quantitative comparison of skeleton-based diffusion on MTT [47]. We compute metrics following STMC [47]. ‘AGD’ denotes the adaptive gradient descent
@@ Appendix C Ablations on Synergistic Energy Fusion
We show the effect of hyper-parameters $\lambda_{l}$, $\lambda_{s}$, and $\lambda_{m}$ (c.f., Equation 10) in Table 7. The results in the first three rows correspond to “Ours (latent only)”, “Ours (semantic only)”, and “Ours” in Table 3, respectively.
Then we conduct ablative experiments on the weights of the two spectra of the energy-based model (latent-aware and semantic-aware), as shown in the middle five rows (Table 7). We find that as the weight of $\lambda_{s}$ increases, the results align more closely with the text (R-Precision and TMR-Score), while larger weights for $\lambda_{l}$ produce smoother motions (Transition distance). This further validates our conclusions in the main paper (Section 5.5 Q2).
@@ Appendix E Ablation Study of Hyper-parameters in Cross-Attention
We investigate the impact of $\gamma_{attn}$ and $\gamma_{reg}$ (c.f., Equation 4. we follow [43] to split $\gamma$ into attention step size $\gamma_{attn}$ and regularization step size $\gamma_{reg}$ for compositional and multi-concept motion generation) in energy-based cross-attention, and the results are presented in Table 9. We notice that $\gamma_{attn},\gamma_{reg}>=0.1$ significantly degrades the performance, and $\gamma_{attn},\gamma_{reg}=[0.001,0.002]$ achieves best results on MTT (Section 5.5 Q1).
@@
Appendix F Energy Distribution Visualization
We show additional contour maps of energy distributions in Figure 6. We provide two examples of the concept conjunction. We visualize the energy distributions of motion latent representations generated from the denoising autoencoder. Multi-concept generation combines concepts from both (a) and (b). Compared with multi-concept generation (i.e., (d) in Figure 6), energy distributions of composed motion (i.e., (c) in Figure 6) show consistent high- and low-energy regions. This demonstrates that complex motion latent distributions can be composed of a set of simple distributions, which also indicates that our method is explainable. Such results further explain the effectiveness of EnergyMoGen (Section 5.5 Q3).
Figure 6: Visual results of energy distributions. For a clear illustration, energy distributions are calculated with interpolation and Gaussian smoothing and then visualized as contour maps. (a) Concept 1, (b) Concept 2, (c) Compositional motion generation, (d) Multi-concept motion generation. Similar regions are highlighted in red
Appendix G More Details on Datasets and Evaluation Metrics
G.1 Datasets
@@
The three datasets use the same motion representation proposed in HumanML3D [19].
G.2 Evaluation Metrics
@@
$\text{FID}=||\mu_{m}-\mu_{\hat{m}}||^{2}-\texttt{TR}(\Sigma_{m}+\Sigma_{\hat{m}}-2\Sigma_{m}\Sigma_{\hat{m}})$ (12)
where $\mu_{gt}$ and $\mu_{\hat{m}}$ are mean of two sets of features. $\Sigma$ is the covariance matrix.
@@
where $N$ is the total number of the motion.
@@
For compositional motion generation, we follow STMC [47] to use R-Precision, TMR-Score, FID, and transition distance to evaluate the performance. Similar to MM-Dist, TMR-Score computes the cosine similarity between the generated motion embedding and text embedding with TMR model [46]. TEACH [2] calculates Euclidean distance between two consecutive frames as transition distance.

### v2

@@ 1 Introduction
In this work, we are interested in the motion latent diffusion model [55, 68], which is another important formulation of diffusion models operating in continuous latent space. However, such models are not feasible for composing motion through algorithms proposed in skeleton-based diffusion models. As latent diffusion models usually use one (or a fixed number of) latent vector(s) to represent variable-length motions (e.g., ranging from 40 to 196 frames), the following issues naturally arise: i) Lack of explicit correspondence (e.g., one-to-one or one-to-many mapping) between latent vectors and motions; ii) It is challenging for latent vectors to support per-frame composition.
@@
Taking these innovations together, EnergyMoGen becomes a general and flexible framework for compositional motion generation (our approach can easily generalize to skeleton-based diffusion models, and the results are provided in the Appendix B). Empirically, we show impressive results on three datasets across three tasks, i.e., HumanML3D [19] and KIT-ML [52] for the text-to-motion generation task, MTT [50] for compositional motion generation and multi-concept motion generation tasks. We conduct comprehensive experiments for in-depth dissection to reveal the effectiveness of EnergyMoGen. Additionally, we introduce a CompML dataset comprising 5000 compositionally generated motion sequences with textual descriptions. We show that the model’s performance can be improved via training on composed motions.
In summary, our contributions are as follows: (1) Our work is the first attempt to solve the motion composition problem from an energy perspective; (2) We explore two spectrums of energy-based models and propose Synergistic Energy Fusion to address critical issues, e.g., text-misalignment and foot sliding; (3) We introduce two critical operations: negation and the combination of conjunction and negation; (4) We generate a CompML dataset via our motion composition method. We demonstrate that training on CompML can improve performance.
@@ 2 Preliminaries: Energy-Based Models
In contrast to explicitly modeling the data distribution, Energy-based models (EBMs) [36, 78, 41] implicitly define the underlying probability density of the distribution through an energy function. For a data sample ${\bm{X}}$, the probability density can be defined via the Boltzmann distribution:
@@
In addition, the modern Hopfield networks [53, 25, 27, 26] interpret the self-attention as EBMs. [53] introduces an energy function $E({\bm{X}},\xi)=-\alpha^{-1}\log\sum^{N}_{i=1}\exp(\alpha({\bm{X}}^{\top}\xi)_{i})+\frac{1}{2}\xi^{\top}\xi$ and a update rule $\xi_{new}=\xi-\gamma(\xi-{\bm{X}}\mathrm{softmax}(\alpha{\bm{X}}^{\top}\xi))$, where the data sample ${\bm{X}}$ consists of $N$ stored patterns (key) ${\bm{x}}_{i}\in\mathbb{R}^{d}$, $\xi\in\mathbb{R}^{d}$ denotes the state vector (query), and $\gamma$ is the update step size. By linearly projecting ${\bm{X}}$ and $\xi$ into Query (${\bm{Q}}$), Key(${\bm{K}}$), and Value(${\bm{V}}$), the attention mechanism can be understood as the energy update ${\bm{Q}}_{new}=\mathrm{softmax}(\alpha{\bm{Q}}{\bm{K}}){\bm{V}}$, when $\gamma=1$ and $\alpha=1/\sqrt{d}$. On this basis, Park et al. [46] extends the theory to cross-attention to handle several tasks, e.g., image editing and inpainting.
@@ 3.1 EnergyMoGen for Motion Generation
We build the architecture of EnergyMoGen upon the motion latent diffusion model [68], which comprises two transformer-based modules: Motion Variational Autoencoder and Motion Latent Diffusion (in Figure 2 (a), (b)).
@@
Building upon the architecture in MLD [68], we propose to employ cross-attention to incorporate text embeddings and motion latent features. In light of recent endorsers on the energy-based attention mechanism [53, 25], we interpret the cross-attention as an energy-based operation to facilitate multi-concept motion generation. Specifically, ${\bm{K}}$ and ${\bm{V}}$ in cross attention is computed as ${\bm{K}}={\bm{c}}W_{\bm{K}}$ and ${\bm{V}}={\bm{c}}W_{\bm{V}}$, where $W_{\bm{K}},W_{\bm{V}}\in\mathbb{R}^{d\times d}$. To alleviate the impact of misalignment in multi-concept text embeddings, we focus on refining ${\bm{c}}$ through adaptive gradient descent (AGD) based on MAP estimation with $p({\bm{K}}|{\bm{Q}})\propto p({\bm{Q}}|{\bm{K}})p({\bm{K}})$. We introduce two energy functions, $E({\bm{K}})$ and $E({\bm{Q}}|{\bm{K}})$, done in the prior work [46]. Then, the gradient of the log distribution can be written as:
@@ 4 Related Works
Compositional Generation. Compositional generation has emerged as a powerful paradigm for creating content (e.g., images and motions) that combines multiple components. Our work is related to certain works in compositional image generation. Some approaches aim to develop cross-attention merging algorithms [12, 46], training strategies [37, 5, 28], and classifier guidance [57, 15], to facilitate this task. Another group of works proposes to use EBMs to compose pre-trained generative models [10, 7, 39, 40, 46, 8]. For example, Du et al. [8] proposed to combine independent trained energy models with several logical operators. Nie et al. [45] and Du et al. [11] introduced Ordinary Different Equation (ODE) and Markov Chain Monte Carlo (MCMC) to optimize the sampling process, respectively. In the field of motion generation, TEACH [2] and SINC [3] are pioneering works for spatial/temporal action composition based on VAE. Subsequent solutions [56, 32, 66] manipulate diffusion models to achieve more natural motion and controllability. However, they modify the reverse process in the skeleton space, which is infeasible for adapting to the latent space.
@@ 5 Experiments
We evaluate our approach on three tasks following prior studies [19, 50]: text-to-motion generation (§$\S$ 5.3), compositional motion generation (§$\S$ 5.4), and multi-concept motion generation (§$\S$ 5.4). Furthermore, we offer an in-depth dissection in §$\S$ 5.5. Please kindly note that our approach can be easily generalized to skeleton-based methods. More information on datasets, evaluation metrics, training details, and the results of skeleton-based diffusion models, can be found in the Appendix A.
@@ 5.5 In-Depth Dissection of EnergyMoGen
Q1: Does the adaptive gradient descent alleviate the issue of text inconsistency in multi-concept motion generation? Yes. As shown in Table 3, our approach with adaptive updating of the textual embedding aligns better with textual descriptions. It improves the R-Precision by 1.3% and 0.9% for Top1 and Top3 accuracy, respectively. Meanwhile, we achieve better FID and comparable results on the TMR-Score and transition distance. The ablation study of $[\gamma_{attn},\gamma_{reg}]$ is provided in the appendix.
Q2: What roles do latent-aware and semantic-aware EBMs play in motion composition? To answer this question, we conduct ablative experiments on MTT [50], results are shown in Table 3. One can figure out that the semantic-aware EBM significantly enhances the text-motion consistency, but it suffers from foot slides and motion jitter. Although the latent-aware method performs poorly on R-Precision and TMR Score, it clearly decreases the transition distance. We notice that Synergistic Energy Fusion effectively alleviates these issues by leveraging the complementary strengths of the two spectrums. We present ablation studies on $\lambda_{s}$, $\lambda_{l}$, and $\lambda_{m}$ in the appendix.
@@ 6 Conclusion
Acknowledgments
This work is supported by the National Natural Science Foundation of China (62472381), Fundamental Research Funds for the Zhejiang Provincial Universities (226-2024-00208), and the Earth System Big Data Platform of the School of Earth Sciences, Zhejiang University.
@@ Acknowledgments
  • Section F: Results of inference time.
  • Section G: Evaluation of foot sliding.
  • Section H: More visual results of energy distributions.
  • Section I: More details on datasets and evaluation metrics.
  • Section J: Limitation and failure case
@@ Appendix A Implementation Details
For the skeleton-based approach, we use an 8-layer transformer with a dimension of 512. We follow [73] to train the model using the Adam optimizer with a batch size of 1024. We train 8000 epochs in total and employ the CosineAnnealing learning policy with the learning rate from 0.0002 to 0.00002.
@@ B.1 Text-to-Motion Generation
We conduct experiments on HumanML3D [19] to evaluate the performance of text-to-motion generation. We use evaluation models from Guo et al. [19] and use the same metrics. The training details of skeleton-based EnergyMoGen are provided in Section A. Experimental results are shown in Table 4. Our approach outperforms current state-of-the-art skeleton-based methods, i.e., ReMoDiffuse [74] and FineMoGen [75] on R-Precision, Diversity, and MM-Dist, while achieving comparable results on FID and MModality.
@@ B.3 Multi-Concept Motion Generation
Table 6: Quantitative comparison of skeleton-based diffusion on MTT [50]. We compute metrics following STMC [50]. ‘AGD’ denotes the adaptive gradient descent.
@@
Figure 6: Ablation of hyper-parameters in Synergistic Energy Fusion on MTT [50].
@@ Appendix C Ablations on Synergistic Energy Fusion
We show the effect of hyper-parameters $\lambda_{l}$, $\lambda_{s}$, and $\lambda_{m}$ in Table 7. The results in the first three rows correspond to “Ours (latent only)”, “Ours (semantic only)”, and “Ours” in Table 3 of the main paper, respectively.
Then we conduct ablative experiments on the weights of the two spectra of the energy-based model (latent-aware and semantic-aware), as shown in the middle five rows (Table 7). We find that as the weight of $\lambda_{s}$ increases, the results align more closely with the text (R-Precision and TMR-Score), while larger weights for $\lambda_{l}$ produce smoother motions (Transition distance).
@@
Note that a more intuitive comparison can be found in Figure 6.
@@ Appendix E Ablation Study of Hyper-parameters in Cross-Attention
We investigate the impact of $\gamma_{attn}$ and $\gamma_{reg}$ We follow [46] to split $\gamma$ into attention step size $\gamma_{attn}$ and regularization step size $\gamma_{reg}$ for compositional and multi-concept motion generation in energy-based cross-attention, and the results are presented in Table 9. We notice that $\gamma_{attn},\gamma_{reg}>=0.1$ significantly degrades the performance, and $\gamma_{attn},\gamma_{reg}=[0.001,0.002]$ achieves best results on MTT.
@@
Appendix F Inference Time
Since our method, like most others, is based on Transformer, we compare its inference time with SOTA Transformer-based diffusion models in Table 10.
Methods FineMoGen MLD GUESS EnergyMoGen
AIT (s) 2.54 0.21 1.79 0.66
Table 10: Inference time. AIT (s) denotes the Average inference time per sentence in seconds.
Appendix G Evaluation of Foot Sliding
Physical Foot Contact score (PFC), proposed in EDGE [62], is used to evaluate the foot sliding problem. We provide a PFC comparison on MTT in Table 11, demonstrating the effectiveness of the proposed Synergistic Energy Fusion.
Methods multi-concept latent-only semantic-only Ours SEF
PFC $\downarrow$ 0.61 0.54 1.05 0.51
Table 11: Evaluation of Foot Sliding. ‘PFC’ denotes the Physical Foot Contact score.
Appendix H Energy Distribution Visualization
We show additional contour maps of energy distributions in Figure 7. We provide two examples of the concept conjunction. We visualize the energy distributions of motion latent representations generated from the denoising autoencoder. Multi-concept generation combines concepts from both (a) and (b). Compared with multi-concept generation (i.e., (d) in Figure 7), energy distributions of composed motion (i.e., (c) in Figure 7) show consistent high- and low-energy regions. This demonstrates that complex motion latent distributions can be composed of simple distributions, indicating that our method is explainable. Such results further explain the effectiveness of EnergyMoGen.
Figure 7: Visual results of energy distributions. For a clear illustration, energy distributions are calculated with interpolation and Gaussian smoothing and then visualized as contour maps. (a) Concept 1, (b) Concept 2, (c) Compositional motion generation, (d) Multi-concept motion generation. Similar regions are highlighted in red.
Appendix I More Details on Datasets and Evaluation Metrics
I.1 Datasets
@@
The three datasets use the same motion representation proposed in [19].
I.2 Evaluation Metrics
@@
$\text{FID}=||\mu_{m}-\mu_{\hat{m}}||^{2}-\texttt{TR}(\Sigma_{m}+\Sigma_{\hat{m}}-2(\Sigma_{m}\Sigma_{\hat{m}})^{\frac{1}{2}})$ (12)
where $\mu_{m}$ and $\mu_{\hat{m}}$ are mean of two sets of features. $\Sigma$ is the covariance matrix.
@@
where $N$ is the total number of motions.
@@
For compositional motion generation, we follow STMC [50] to use R-Precision, TMR-Score, FID, and transition distance to evaluate the performance. Similar to MM-Dist, TMR-Score computes the cosine similarity between the generated motion embedding and text embedding with the TMR model [49]. TEACH [2] calculates the Euclidean distance between two consecutive frames as the transition distance.
Figure 8: Failure case.
Appendix J Limitation and Failure Case
Existing latent diffusion models encode motion into a single (or a fixed number of) latent vector(s), which limits the use of per-frame composition algorithms. We propose using an energy function to directly model the latent vector(s) encapsulating the overall features, e.g., temporal and skeletal features. The energy function (or energy) is additive. This property enables motion composition by composing energy functions (generated latent vectors) from different concepts together via conjunction and negation. However, our method still struggles with completely novel concepts. We provide a failure case in Figure 8.